Can someone please help me with the work for this problem: 6x^2-11x-10 I need to factor it completely and I am having trouble without a common factor. Thank you so much in advance!
Answer by checkley71(8403)

6X~2-11X-10=0 OR (3X+2)(2X-5)=0 OR 3X+2=0 OR 3X=-2 OR X=-2/3 & 2X-5=0 OR 2X=5 OR X=5/2 OR X=2.5